# Ambient Noise Full Waveform Inversion with Neural Operators

### v2

@@ 3.2 Data-Driven Training
In this study, we train a single 2D HNO for linear nodal arrays across the northern LA basins. The model is trained in a supervised learning manner and with synthetic data generated from a spectral element method (SEM) using the software SALVUS (Afanasiev et al., 2019). The computational domain is set to $80$ km (horizontal) $\times$ $40$ km (vertical) on a $256\times 128$ mesh, which accommodates the longest seismic line in the area. We simulate $40$-s-duration wavefields with a time step of $0.001$ s, which are excited by a vertical force and Ricker wavelet time function with a central frequency of $0.3$ Hz. For ambient noise data applications, the source is randomly placed along the free surface from a uniform distribution of horizontal position. The $V_{P}$ and $V_{S}$ models are generated from random fields, the parameters of which are provided in Table S2. The density models are derived from $V_{P}$ using the empirical relation by Brocher (2005), which are input to SALVUS but not to the HNO. The time-domain solutions from the SEM solver are Fourier-transformed and filtered to the frequency band of interest ($0.1$ to $0.5$ Hz) for use with the HNO. We train an HNO using $27000$ simulations and validate the model performance using $3000$ simulations (where a simulation means a single source and a particular velocity model). We define the loss function as
$Loss=0.95\frac{\left\|\mathbf{u}-\mathbf{\hat{u}}\right\|_{1}}{\left\|\mathbf{u}\right\|_{1}}+0.05\frac{\left\|\mathbf{u}-\mathbf{\hat{u}}\right\|_{2}}{\left\|\mathbf{u}\right\|_{2}},$ (15)
where $\mathbf{u}$ and $\mathbf{\hat{u}}$ denote the true and predicted wavefields, respectively. We use an Adam optimizer (Kingma & Ba, 2014) with a learning rate of $0.001$ and a scheduler that decays the learning rate by half every $30$ epochs. The batch size is $256$. Figure S1 shows the loss curves. The full training process took $2$ days with $8$ NVIDIA RTX A6000 GPUs. Note that the training is a one-time effort and no further training is required for different seismic lines or the inversion stage. Table 1 compares the time consumed by SALVUS and the trained HNO in forward modeling and FWI under the same configuration. In performing FWI for this 2D case, the HNO with AD is approximately two orders of magnitude faster than the SEM with the adjoint method. The computational advantage will only be greater in 3D (Zou et al., 2024).
Table 1: Comparison of time consumed by SALVUS and HNO in forward modeling and FWI. The experiments are conducted on a community model called CVM-S4.26 using a single NVIDIA RTX A6000 GPU. The forward modeling time is measured for a single event with one source and $234$ receivers on the free surface. The FWI time is measured for one tomographic iteration for the same event configuration.
@@ 3.3 Synthetic Tests
To test generalization, we first evaluate the HNO performance with the synthetic data generated from a community velocity model named CVM-S4.26 (Lee et al., 2014). We extract a profile of the 3D CVM-S4.26 along the longest seismic line, SB1, in the BASIN survey. Figure 2 shows the HNO-predicted results for a sample shot, with the SALVUS-simulated waveforms serving as the ground truth. The direct output of the HNO is in the frequency domain and is Fourier-transformed to the time domain for visualization purposes. The forward prediction by the HNO is not perfect with a cross-correlation coefficient of 0.98, but we will show that it is sufficiently accurate in the upcoming FWI experiment.
@@
Before the inversion, we compute the sensitivity kernels for the HNO-AD method (Equation 14), following the definition in Tromp et al. (2005). In Figure 3, we put sources and receivers at nodes on the free surface and sum the gradients of the misfit with respect to the velocity parameters over all source-receiver pairs. The misfit is defined as the squared error between the HNO-predicted wavefield with a 1D initial model and the SALVUS-simulated wavefield with the true model (CVM-S4.26). We average the sensitivity kernels horizontally to obtain functions of depth. It can be seen that the sensitivity drops dramatically with depth, which indicates that updates to deeper parts of the velocity models are less reliable. Therefore, we will only present the inversion results for the top $10$ km hereafter.
@@
Figure 3: Sensitivity kernels for the HNO-AD method, defined as the gradients of the misfit with respect to the velocity parameters, which are summed over all source-receiver pairs and averaged horizontally. The misfit is defined as the squared error between the HNO-predicted wavefield with a 1D initial model and the SALVUS-simulated wavefield with the true model (CVM-S4.26). The sensitivity values for $V_{P}$ and $V_{S}$ are each normalized by its maximum amplitude.
@@
In the inversion stage, we freeze the HNO parameters and update $V_{P}$ and $V_{S}$ simultaneously with gradients calculated from AD. The inversion is also performed in the frequency domain, and the objective function is defined as the mean squared error (MSE) between the HNO-predicted and SALVUS-simulated wavefields. For use with ambient noise data where the amplitude is often not well-defined, we normalize the amplitude to unity before computing the MSE. We add no regularization term to the objective function but implicitly impose some regularization by smoothing the gradients with a Gaussian filter (Tape et al., 2010; Zhu et al., 2015). The smoothing radius is set to 3×3$\times$3 × grid spacing ($945$ m). We average the CVM-S4.26 profile horizontally to get a 1D initial model. We start with lower frequency data and gradually feed in higher frequency components, which is also common practice in traditional FWI (Wang et al., 2018; Zhang et al., 2018). This is because lower frequency data are usually less noisy, easier to fit, and better for global convergence. Also, the neural operator performs more accurately for lower frequencies (Zou et al., 2024).
We also use an Adam optimizer for gradient descent in FWI, with a learning rate of $0.05$ and a scheduler that decays the learning rate by half every $5$ epochs. The gradients are summed over a mini-batch of $64$ data points, where a single data point corresponds to a single frequency and a single source. We monitor MSE for the full frequency band while progressively incorporating higher frequency data as inversion proceeds, as shown in Figure 4. We conduct FWI experiments on both noise-free data and data perturbed with different levels of Gaussian random noise. We activate $234$ sources along the free surface of the CVM-S4.26 slice for SB1, located at the nodes. The receivers are also placed at the nodes.
@@
Figure 4: Overall MSE (mean squared error) in the FWI process, where higher frequency data is progressively incorporated. We use an Adam optimizer with a learning rate of $0.05$ and a scheduler that decays the learning rate by half every $5$ epochs. The batch size is $64$.
@@
Figure 5 shows a sample shot of synthetic data perturbed with different levels of noise. The noise is generated from a zero-mean Gaussian distribution with a standard deviation (SD) equal to a factor times the SD of the noise-free data. The corresponding FWI results for noise-free data and noisy data are displayed in Figure 6. The noise-free inversion in the second row clearly reveals the basin model, demonstrating the forward modeling accuracy of the HNO from another perspective. The deeper part of the model is not well resolved because the source-receiver geometry limits sensitivity to near-surface regions. When 5×5$\times$5 × noise is added, the FWI results maintain good quality, even though the seismic signal already becomes unrecognizable to the eye. We increase the noise level until the inversion is on the verge of collapse, with a perturbation of 10×10$\times$10 × noise that totally obliterates the waveforms.
@@
Figure 5: A sample shot of synthetic data perturbed with different levels of noise. The noise is generated from a zero-mean Gaussian distribution with a SD equal to a factor times the SD of the noise-free data.
@@
These synthetic tests serve to verify the generalizability of the neural operator trained with random fields and noise-free data to realistic velocity structures and noisy data, supervised by some definition of ground truth that is unavailable in real data. They show the excellent performance of the HNO-AD method in low signal-to-noise ratio (SNR) situations. Additionally, the entire FWI process can be completed in minutes. In contrast, FWI with the adjoint-state method takes $16$ hours for $10$ iterations and makes little progress with the same 1D initialization even for noise-free data (Figure S2).
@@ 3.4 Application to Real Seismic Data from the Broader Los Angeles Region
We apply the trained neural operator to a real-world ambient noise tomography example. The real data comes from the BASIN survey that deployed $10$ linear nodal arrays in the San Gabriel (SG), Chino, and San Bernardino (SB) basins north of LA (Clayton et al., 2019). $758$ Fairfield ZLand three-component nodes with a corner frequency of $5$ Hz recorded ambient noise for approximately a month. The spacing ranges from $250$ to $300$ m. Figure 7 shows a map of the study area and station configuration. The sedimentary formations date from the opening of the LA basin in the Miocene (Wright, 1991). Details in the geologic setting can be found in Villa et al. (2023) and Ghose et al. (2023).
@@
where $G_{AB}(t)$ is the displacement EGF between two stations $A$ and $B$, and $C_{AB}(t)$ is the cross-correlation between the seismograms recorded at these two stations. The positive and negative lags of $C_{AB}$ are averaged to enhance the SNR and to mitigate the effect of inhomogeneous source distribution (Lin et al., 2008). Our cross-correlation workflow is based on Bensen et al. (2007). Equation 16 is indeed a far-field approximation, which holds for inter-station distance greater than three wavelengths (Yao et al., 2006). However, in practice, this condition is often relaxed to make better use of the available data (Li et al., 2023; Wu et al., 2021; Zhang et al., 2018). Especially for a local-scale survey like BASIN, we set the minimum inter-station distance to $5$ km for use with $0.2$ to $0.5$ Hz data. Increasing this distance to $10$ km yields a very similar result (Figure S3).
The EGF obtained from cross-correlation accounts for wave physics in a 3D medium, to which a phase shift of $\pi/4$ must be applied prior to a 2D inversion (Forbriger et al., 2014; Schäfer et al., 2014; Zhang et al., 2018). Figure S4 shows the effect of this transformation. A factor for amplitude correction is further needed, but we only use the phase information, as the amplitude recovered in ANT is not well-defined due to various normalizations done in data processing (Bensen et al., 2007). We have verified through synthetic tests that the phase information is sufficient to reconstruct the velocity models (Figure 6). We focus solely on the phase information by normalizing the amplitude to unity in inversion. This helps to bridge the gap between the synthetic Green’s function (SGF) and EGF, as they have different source bandwidths. We match the Z-Z and Z-R cross-correlations with the vertical and horizontal wavefield components predicted by the HNO, respectively. While we mainly retrieve Rayleigh waves with particle motion confined in a 2D plane, we will also discuss the use of Love waves through T-T correlations for the 2D anti-plane problem in the next section. In this study, we follow the principle of Green’s function retrieval, while Tromp et al. (2010), Sager et al. (2018, 2020), and Tsai et al. (2024) established the full waveform ambient noise inversion technique that goes beyond that by interpreting correlation functions as self-consistent observables.
@@
where $M$ is the number of grid points on which the velocity models are discretized.
@@ 4.1 San Gabriel and Chino Basins
We begin the real data FWI with the longest line, SB1, which extends about $80$ km from west to east. Beneath SB1 lie two significant basins — the deeper SG basin in the west and the shallower Chino basin in the east. The depth of the sediment-basement interface has been constrained by Villa et al. (2023). In addition to this 3D basin depth model and the CVM-S4.26 (Lee et al., 2014), another reference is from Li et al. (2023) who constructed a 3D shear wave velocity model for the same area through surface wave dispersion analysis. We start the inversion with a 1D model calculated from this reference model, which results in Figure 8a. Overall, the inverted velocity models reveal basin structures that align well with Villa et al. (2023), capturing significant features including the deeper and more heterogeneous SG basin in the west and the low-velocity zone at the edge of the relatively flat Chino basin adjacent to the SB basin. We demonstrate the robustness of this result through another two initial models, the smoothed models from Li et al. (2023) (Figure 8b) and CVM-S4.26 (Figure 8c). Results from these different initial models show excellent overall consistency, although they differ in details. In particular, the 1D initial model can achieve performance comparable to that of models with more detailed priors.
@@
Figure 8: Inversion results for SB1 with different initial models including the (a) 1D and (b) smoothed models from Li et al. (2023), and (c) the smoothed CVM-S4.26. The along-profile distance increases from west to east. The basin bottom from Villa et al. (2023) is delineated with black dashed lines for reference.
@@
The incorporation of higher frequency data in later stages of FWI contributes to refining the shallower structure but also introduces high-frequency noise. Meanwhile, the inaccuracy of the neural operator in forward modeling introduces another source of uncertainty. We compare SALVUS-simulated waveforms (SGFs) using the CVM-S4.26, model from Li et al. (2023), and our model from Figure 8a with the EGF for SB1 in Figure 9. The corresponding waveform difference is plotted in Figure S5. The point here is not to conclude that our model is better than others, as we have not incorporated the full wave physics in 3D, but rather to demonstrate that the HNO-AD method can achieve accuracy comparable to traditional methods at a much lower cost in terms of time and effort. Also note that our model starts with a 1D initial model, while the other two models have stronger priors (Lee et al., 2014; Li et al., 2023).
@@
Figure 9: Examples of SALVUS-simulated waveforms (SGFs) using the CVM-S4.26, model from Li et al. (2023), and our model from Figure 8a compared with the EGF for SB1. Only $V_{S}$ is plotted for simplicity. The frequency band of the data is $0.2$ to $0.5$ Hz. In plotting, the waveform is normalized by trace and every third trace is displayed.
@@
The same neural operator can be directly applied to any seismic line in the study region. We show another two examples for SB4 and SG1 in Figures 10 and 11, respectively. For SB4 that crossed the Chino basin from north to south, our inversion result reveals a low-velocity zone at the southern end near the Chino fault (Figure 7), which is also seen in the CVM-S4.26 and Li et al. (2023). The low-velocity anomalies between 00 and $15$ km present in Li et al. (2023) are absent in both our model and CVM-S4.26. While the authors did not provide specific interpretation, the seismic line crossed the Red Hill fault at approximately $5$ km, which could explain the observed discontinuity (if it is not an artifact). Interestingly, our model shows a velocity jump near the Red Hill fault. Overall, the tomography for SB4 displays a shallow (approximately $2$ km) sedimentary basin in agreement with previous studies. The sediment-basement interface beneath SG1 is much deeper, as captured in all the displayed models. Our result for SG1 aligns with that for SB1 in the SG basin in terms of basin depth but captures less heterogeneity, which might be limited by the array length. These plug-and-play inversions can be completed in a few minutes.
@@
Figure 10: Inversion results for SB4, in comparison to the CVM-S4.26 and model from Li et al. (2023). Only $V_{S}$ is plotted for simplicity. The along-profile distance increases from north to south. The basin bottom from Villa et al. (2023) is delineated with black dashed lines for reference.
@@
Figure 11: Inversion results for SG1, in comparison to the CVM-S4.26 and model from Li et al. (2023). Only $V_{S}$ is plotted for simplicity. The along-profile distance increases from north to south. The basin bottom from Villa et al. (2023) is delineated with black dashed lines for reference.
4.2 Use of Love Waves from T-T Correlations
In previous sections, we focused on the 2D in-plane problem (Equation 1) associated with Rayleigh waves and inverted for $V_{P}$ and $V_{SV}$ models. The $V_{SH}$ model can be separately derived from Love waves with particle motion perpendicular to the plane through T-T cross-correlations $u_{T}$, which is termed the anti-plane problem:
$\frac{1}{\rho V_{SH}^{2}}\frac{\partial^{2}u_{T}}{\partial t^{2}}-\nabla\cdot\left(\frac{1}{\rho}\nabla u_{T}\right)=f.$ (18)
In this section, we invert for the $V_{SH}$ model for SB1 following the same methodology to complete the 2D FWI. $V_{SV}=V_{SH}$ only holds when the medium is isotropic. We define the radial anisotropy as in Jaxybulatov et al. (2014):
$\xi=\frac{2(V_{SH}-V_{SV})}{V_{SH}+V_{SV}}.$ (19)
Figure 12 compares the inverted $V_{SV}$ and $V_{SH}$ models for SB1 and plots the radial anisotropy. The same 1D initial model as in Figure 8a is used for both cases and we take the relatively robust results from low frequencies ($0.2$ to $0.3$ Hz) for comparison. The overall feature of the basin depth variation is captured in both inversions, while $V_{SV}$ seems to penetrate deeper and $V_{SH}$ provides better constraint for the shallower part. The radial anisotropy shows a clear sign flip along an interface that is likely to indicate changes in lithology. The near-surface negative anisotropy can result from vertical cracks caused by repeated earthquakes, which becomes positive with increasing depth and pressure due to the horizontal layering of sedimentary rocks (Jiang & Denolle, 2022). We are less convinced by the magnitude and depth resolution of the anisotropy and the reliability decreases with depth due to loss in sensitivity. It is important to note that the primary goal of this experiment, and indeed the entire study, is to demonstrate the functionality of a new, efficient, and generalized method rather than to delve into a detailed interpretation of any specific area. This is also why we chose a well-studied area, ensuring that we have results for validation.
Figure 12: Inverted $V_{SV}$ and $V_{SH}$ for SB1 and the radial anisotropy, using data from $0.2$ to $0.3$ Hz. The same 1D initial model as in Figure 8a is used for both cases.

### v3

@@ 2.2 Data-Driven Training
In this study, we train a single 2D HNO for linear nodal arrays across the northern LA basins. The model is trained in a supervised learning manner using synthetic data generated from a spectral element method (SEM) using the software SALVUS (Afanasiev et al., 2019). The computational domain is set to 808080 km (horizontal) $\times$ 404040 km (vertical) on a 256×128256$\times$ 128256 × 128 mesh, which accommodates the longest seismic line, SB1, in the area. We simulate 404040-s-duration wavefields with a time step of 0.0010.0010.001 s, which are excited by a vertical force and Ricker wavelet time function with a central frequency of 0.30.30.3 Hz. For ambient noise data applications, the source is randomly placed along the free surface from a uniform distribution of horizontal position. The $V_{P}$ and $V_{S}$ models are generated from random fields sampled around a background 1D model, which is averaged along the SB1 line from a reference model named CVM-S4.26 (Lee et al., 2014). The parameters for the random fields are provided in Table S2. The density models are derived from $V_{P}$ using the empirical relation by Brocher (2005), which are input to SALVUS but not to the HNO. The time-domain solutions from the SEM solver are Fourier-transformed and filtered to the frequency band of interest (0.10.10.1 to 0.50.50.5 Hz) for use with the HNO. We train an HNO using 270002700027000 simulations and validate the model performance using 300030003000 simulations (where a simulation means a single source and a particular velocity model). We define the loss function as a combination of relative $\ell_{1}$-norm and $\ell_{2}$-norm of the data residual:
[…]
where $\mathbf{d}$ and $\mathbf{\hat{d}}$ denote the true and predicted data, respectively. We use an Adam optimizer with a learning rate of 0.0010.0010.001 and a scheduler that decays the learning rate by half every 303030 epochs. The batch size is 256256256. Figure S1 shows the loss curves. The full training process took 222 days with 888 NVIDIA RTX A6000 GPUs. Note that the training is a one-time effort and no further training is required for different seismic lines or the inversion stage, granted the data have a similar scale. Table 1 compares the time consumed by SALVUS and the trained HNO in forward modeling and FWI under the same configuration. In performing FWI for this 2D case, the HNO with AD is approximately two orders of magnitude faster than the SEM with the adjoint method. The computational advantage will only be greater in 3D (Zou et al., 2024).
Table 1: Comparison of time consumed by SALVUS and HNO in forward modeling and FWI. The experiments are conducted on a community model called CVM-S4.26 using a single NVIDIA RTX A6000 GPU. The forward modeling time is measured for a single event with one source and 234234234 receivers on the free surface. The FWI time is measured for one tomographic iteration for the same event configuration.
@@ 2.3 Synthetic Tests
To test generalization, we first evaluate the HNO performance with the synthetic data generated from a community velocity model named CVM-S4.26 (Lee et al., 2014). We extract a profile of the 3D CVM-S4.26 along the longest seismic line, SB1, in the BASIN survey. This model is never explicitly seen by the HNO. Figure 2 shows the HNO-predicted results for a sample shot excited by a vertical force and Ricker wavelet as in training, with the SALVUS-simulated waveforms serving as the ground truth. The direct output of the HNO is in the frequency domain (0.10.10.1 to 0.50.50.5 Hz) and is inverse Fourier-transformed to the time domain for visualization purposes. The forward prediction by the HNO is not perfect with a cross-correlation coefficient of 0.98, but we will show that it is sufficiently accurate in the upcoming FWI experiment.
@@
Before the inversion, we compute the sensitivity kernels for the HNO-AD method, following the definition in Tromp et al. (2005). In Figure 3, we put sources and receivers at nodes on the free surface and sum the gradients of the misfit with respect to the velocity parameters over all source-receiver pairs. The misfit is defined as the squared error between the HNO-predicted data with a 1D initial model and the SALVUS-simulated data with the true model (CVM-S4.26). The 1D initial model is obtained by horizontally averaging the CVM-S4.26 profile. We average the sensitivity kernels horizontally to obtain functions of depth. We can see that the sensitivity drops dramatically with depth, which indicates that updates to deeper parts of the velocity models are less reliable. Therefore, we will only present the inversion results for the top 101010 km hereafter.
@@
Figure 3: Sensitivity kernels for the HNO-AD method, defined as the gradients of the misfit with respect to the velocity parameters, which are summed over all source-receiver pairs and averaged horizontally. The misfit is defined as the squared error between the HNO-predicted data with a 1D initial model and the SALVUS-simulated data with the true model (CVM-S4.26). The sensitivity values for $V_{P}$ and $V_{S}$ are each normalized by its maximum amplitude.
@@
In the inversion stage, we freeze the HNO parameters and update $V_{P}$ and $V_{S}$ simultaneously with gradients calculated from AD. The inversion is also performed in the frequency domain, and the objective function is defined as the mean squared error (MSE) between the HNO-predicted and SALVUS-simulated data. For use with ambient noise data where the amplitude is often not well-defined, we normalize the amplitude to unity before computing the MSE. This is done by dividing the frequency-domain data, a complex-valued quantity, by its amplitude. We add no regularization term to the objective function but implicitly impose some regularization by smoothing the gradients with a Gaussian filter (Tape et al., 2010; Zhu et al., 2015). The smoothing radius is set to 3×3$\times$3 × grid spacing (945945945 m) in both horizontal and vertical directions. We average the CVM-S4.26 profile horizontally to get a 1D initial model. We start with lower frequency data and gradually feed in higher frequency components (from 0.10.10.1 to 0.50.50.5 Hz), which is also common practice in traditional FWI to mitigate cycle-skipping (Wang et al., 2018; Zhang et al., 2018). This is because skipping one cycle in lower frequency data requires larger velocity variation, whereas velocity variation is smaller at greater depth — where lower frequency data has better sensitivity. Also, the neural operator performs more accurately for lower frequencies (Zou et al., 2024). Figure 4 shows the MSE for the full frequency band in the inversion process. We use an Adam optimizer with a learning rate of 0.050.050.05 and a scheduler that decays the learning rate by half every 555 epochs. The gradients are summed over a mini-batch of 646464 data points, where a single data point corresponds to a single frequency and a single source.
@@
Figure 4: Overall MSE (mean squared error) in the FWI process, where higher frequency data is progressively incorporated. We use an Adam optimizer with a learning rate of 0.050.050.05 and a scheduler that decays the learning rate by half every 555 epochs. The batch size is 646464.
@@
We conduct FWI experiments on both noise-free data and data perturbed with different levels of Gaussian random noise. We activate 234234234 sources along the free surface of the CVM-S4.26 slice for SB1, located at the nodes. The receivers are also placed at the nodes. Figure 5 shows a sample shot of synthetic data perturbed with different levels of noise. The noise is generated from a zero-mean Gaussian distribution with a standard deviation (SD) equal to a factor times the SD of the noise-free data. The corresponding FWI results for noise-free data and noisy data are displayed in Figure 6. The noise-free inversion in the second row clearly reveals the basin model, demonstrating the forward modeling accuracy of the HNO from another perspective. The deeper part of the model is not well resolved because the source-receiver geometry and frequency content limit sensitivity to near-surface regions. When 5×5$\times$5 × noise is added, the FWI results maintain good quality, even though the seismic signal already becomes unrecognizable to the eye. This could be explained by the fact that FNOs are global operators, and the zero-mean noise might be partially averaged out in the kernel. However, real noise typically does not behave this nicely. We increase the noise level until the inversion is on the verge of collapse, with a perturbation of 10×10$\times$10 × noise that totally obliterates the waveforms.
@@
Figure 5: A sample shot of synthetic data perturbed with different levels of noise. The noise is generated from a zero-mean Gaussian distribution with an SD equal to a factor times the SD of the noise-free data.
@@
These synthetic tests serve to verify the generalizability of the neural operator trained with random fields and noise-free data to realistic velocity structures and noisy data, supervised by some definition of ground truth that is unavailable in real data. They show the excellent performance of the HNO-AD method in low signal-to-noise ratio (SNR) situations. Additionally, the entire FWI process can be completed in minutes. In contrast, FWI with the adjoint-state method takes 161616 hours for 101010 iterations and makes little progress with the same 1D initialization even for noise-free data (Figure S2). We attribute the main difference to the use of mini-batch Adam optimization in the HNO-AD method, which is not yet supported by SALVUS.
@@ 2.4 Application to Real Seismic Data from the Broader Los Angeles Region
We apply the trained neural operator to a real-world ambient noise tomography example. The real data comes from the BASIN survey that deployed 101010 linear nodal arrays in the San Gabriel (SG), Chino, and San Bernardino (SB) basins north of LA (Clayton et al., 2019). A total of 758758758 Fairfield ZLand three-component nodes with a corner frequency of 555 Hz recorded ambient noise for approximately a month. Each line consisted of 141414 to 260260260 geophones, spaced approximately 0.250.250.25 km apart. The geometry of the survey was designed to detect subsurface variations up to 0.50.50.5 km in horizontal scale. Figure 7 shows a map of the study area and station configuration. The sedimentary formations date from the opening of the LA basin in the Miocene (Wright, 1991). Details in the geologic and tectonic setting can be found in Villa et al. (2023) and Ghose et al. (2023).
@@
where $G_{AB}(t)$ is the displacement EGF between two stations $A$ and $B$, and $C_{AB}(t)$ is the cross-correlation between the seismograms recorded at these two stations. The positive and negative lags of $C_{AB}$ are averaged to enhance the SNR and to mitigate the effect of inhomogeneous source distribution (Lin et al., 2008). Although more effective methods for addressing non-diffuse noise fields have been explored (Liu, 2020; Liu et al., 2023a), this is not the focus of this study, and we follow the standard procedure. Our cross-correlation workflow is based on Bensen et al. (2007) and Li et al. (2023b). We correlate both daytime and nighttime recordings from the one-month deployment, since using only nighttime data merely reduced the anthropogenic noise in a minor way (Li et al., 2023b). One-hour data segments are correlated and stacked to derive the final correlation. We convert the three-component data from the vertical-north-east (ZNE) to the vertical-radial-transverse (ZRT) coordinate system. The particle motion of Rayleigh waves lies in the ZR plane, whereas Love waves involve mainly horizontal motion in the transverse direction. We demean the raw seismograms and apply a bandpass filter between 0.10.10.1 and 222 Hz. Prior to the correlation, we perform temporal normalization to suppress the effects of earthquakes and apply spectral whitening to broaden the effective bandwidth.
The EGF obtained from cross-correlation accounts for wave physics in a 3D medium, to which a phase shift of $\pi/4$ must be applied prior to a 2D inversion (Forbriger et al., 2014; Schäfer et al., 2014; Zhang et al., 2018). Figure S3 shows the effect of this transformation. A factor for amplitude correction related to geometrical spreading is further needed (Zhang et al., 2018), but we focus solely on the phase information, as the amplitude recovered in ANT is not well-defined. We have verified through synthetic tests that the phase information is sufficient to reconstruct the velocity models (Figure 6). We eliminate the effect of amplitude differences between 2D and 3D by normalizing the amplitude to unity during inversion. This also helps to bridge the gap between the synthetic Green’s function (SGF) and EGF, as they have different source bandwidths. We match the Z-Z and Z-R cross-correlations with the vertical and horizontal waveforms predicted by the HNO, respectively. We mainly retrieve Rayleigh waves with particle motion confined in a 2D plane. In this study, we follow the principle of Green’s function retrieval, while Tromp et al. (2010), Sager et al. (2018, 2020), and Tsai et al. (2024) established the full waveform ambient noise inversion technique that goes beyond that by interpreting correlation functions as self-consistent observables.
@@
where $M$ is the number of grid points on which the velocity models are discretized. This is to constrain $V_{P}$, which is not well constrained by surface waves.
@@ 3 Results
We begin the real data FWI with the longest line, SB1, which extends about 808080 km from west to east. Beneath SB1 lie two significant basins — the deeper SG basin in the west and the shallower Chino basin in the east. The depth of the sediment-basement interface has been constrained by Villa et al. (2023). In addition to this 3D basin depth model and the CVM-S4.26 (Lee et al., 2014), another reference is from Li et al. (2023b) who constructed a 3D shear wave velocity model for the same area through surface wave dispersion analysis. We start the inversion with a 1D model calculated from this reference model, which results in Figure 8a. Overall, the inverted velocity models reveal basin structures that align well with Villa et al. (2023), capturing significant features including the deeper and more heterogeneous SG basin in the west and the low-velocity zone at the edge of the relatively flat Chino basin adjacent to the SB basin. We demonstrate the robustness of this result through another two initial models, the smoothed models from Li et al. (2023b) (Figure 8b) and CVM-S4.26 (Figure 8c). Results from these different initial models show excellent overall consistency, although they differ in details. In particular, the 1D initial model can achieve performance comparable to that of models with more detailed priors.
@@
Figure 8: Inversion results for SB1 with different initial models including the (a) 1D and (b) smoothed models from Li et al. (2023b), and (c) the smoothed CVM-S4.26. The along-profile distance increases from west to east. The basin bottom from Villa et al. (2023) is delineated with black dashed lines for reference.
@@
The incorporation of higher frequency data in later stages of FWI contributes to refining the shallower structure but also introduces high-frequency noise. Some of the noise can potentially be suppressed using a signal window. Meanwhile, the inaccuracy of the neural operator in forward modeling introduces another source of uncertainty. We compare SALVUS-simulated waveforms (SGFs) using the CVM-S4.26, model from Li et al. (2023b), and our model from Figure 8a with the EGF for SB1 in Figure 9. The corresponding waveform difference is plotted in Figure S4. The point here is not to conclude that our model is better than others, but rather to demonstrate that the HNO-AD method can achieve accuracy comparable to traditional methods at a much lower cost in terms of time and effort. Also note that our model starts with a 1D initial model, while the other two models have stronger priors (Lee et al., 2014; Li et al., 2023b).
@@
Figure 9: Examples of SALVUS-simulated waveforms (SGFs) using the CVM-S4.26, model from Li et al. (2023b), and our model from Figure 8a compared with the EGF for SB1. Only $V_{S}$ is plotted in the top left for simplicity. The frequency band of the data is 0.20.20.2 to 0.50.50.5 Hz. In plotting, the waveform is normalized by trace and every third trace is displayed.
@@
The same neural operator can be directly applied to any seismic line in the study region, because it was trained to cover the longest one. While the model cannot extrapolate to larger domains than it was trained on, application to smaller domains is straightforward by simply extracting a subset of the inversion result. Additionally, no specific orientation is assumed for the 2D profiles, since the training data are generated in a Cartesian coordinate system without reference to geological orientation. This ensures that the model is not biased toward any particular direction. We show another two examples for SB4 and SG1 in Figures 10 and 11, respectively. For SB4 that crossed the Chino basin from north to south, our inversion result reveals a low-velocity zone at the southern end near the Chino fault (Figure 7), which is also seen in the CVM-S4.26 and Li et al. (2023b). The low-velocity anomalies between 0 and 151515 km present in Li et al. (2023b) are absent in both our model and CVM-S4.26. While the authors did not provide specific interpretation, the seismic line crossed the Red Hill fault at approximately 555 km, which could explain the observed discontinuity (if it is not an artifact). Interestingly, our model shows a velocity jump near the Red Hill fault. Overall, the tomography for SB4 displays a shallow (approximately 222 km) sedimentary basin in agreement with previous studies. The sediment-basement interface beneath SG1 is much deeper, as captured in all the displayed models. Our result for SG1 aligns with that for SB1 in the SG basin in terms of basin depth but captures less heterogeneity, which might be limited by the array length. These inversions can be completed in a few minutes.
@@
Figure 10: Inversion results for SB4, in comparison to the CVM-S4.26 and model from Li et al. (2023b). Only $V_{S}$ is plotted for simplicity. The along-profile distance increases from north to south. The basin bottom from Villa et al. (2023) is delineated with black dashed lines for reference.
@@
Figure 11: Inversion results for SG1, in comparison to the CVM-S4.26 and model from Li et al. (2023b). Only $V_{S}$ is plotted for simplicity. The along-profile distance increases from north to south. The basin bottom from Villa et al. (2023) is delineated with black dashed lines for reference.